# Skywork R1V2: Multimodal Hybrid Reinforcement Learning for Reasoning

### v1

@@ Abstract
We present Skywork R1V2, a next-generation multimodal reasoning model and a major leap forward from its predecessor, Skywork R1V. At its core, R1V2 introduces a hybrid reinforcement learning paradigm that harmonizes reward-model guidance with rule-based strategies, thereby addressing the long-standing challenge of balancing sophisticated reasoning capabilities with broad generalization. To further enhance training efficiency, we propose the Selective Sample Buffer (SSB) mechanism, which effectively counters the “Vanishing Advantages” dilemma inherent in Group Relative Policy Optimization (GRPO) by prioritizing high-value samples throughout the optimization process. Notably, we observe that excessive reinforcement signals can induce visual hallucinations—a phenomenon we systematically monitor and mitigate through calibrated reward thresholds throughout the training process. Empirical results affirm the exceptional capability of R1V2, with benchmark-leading performances such as 62.6 on OlympiadBench, 79.0 on AIME2024, 63.6 on LiveCodeBench, and 74.0 on MMMU. These results underscore R1V2’s superiority over existing open-source models and demonstrate significant progress in closing the performance gap with premier proprietary systems, including Gemini 2.5 and OpenAI o4-mini. The Skywork R1V2 model weights have been publicly released to promote openness and reproducibility***https://huggingface.co/Skywork/Skywork-R1V2-38B.
@@ 1 Introduction
Recent advancements in "slow-thinking" multimodal models—such as OpenAI-o1, Gemini-Thinking, Kimi-1.5, and Skywork-R1V—have led to substantial progress in complex reasoning tasks across mathematics and science. These models emulate reflective cognitive processes, achieving stronger performance than their "fast-thinking" counterparts, including OpenAI-o4 and Claude-3.5, which prioritize rapid generation over deliberate reasoning. For instance, slow-thinking models demonstrate over 30% improvement on benchmarks like AIME24 and AMC23, along with approximately 10% gains on science-related benchmarks such as GPQA.
@@
However, extending slow-thinking strategies to multimodal domains introduces new challenges. While improvements are observed on visual reasoning tasks such as MMMU and MathVision, performance degrades on general perception benchmarks like AI2D and OCRBench. This decline is often accompanied by a rise in visual hallucinations, highlighting a critical issue: How can we effectively promote slow-thinking behavior in vision-language models (VLMs) without compromising their generalization capabilities?
@@
To tackle this problem, we introduce Skywork R1V2, a next-generation vision-language model that acquires multimodal reasoning skills directly via reinforcement learning (RL), bypassing the need for teacher model distillation. R1V2 is trained on a comprehensive dataset spanning visual perception, scientific inquiry, and abstract reasoning, encompassing both general-purpose and reasoning-intensive tasks. The training process utilizes hybrid optimization signals derived from the R1V-RM reward model and structured rule-based feedback.
@@
At its core, R1V2 builds upon the Group Relative Policy Optimization (GRPO) algorithm, which computes relative advantages by comparing candidate responses within the same query group. As training progresses, however, responses may converge (i.e., become uniformly correct or incorrect), resulting in vanishing advantage signals and limited reasoning diversity. To address this, we propose a Selective Sample Buffer (SSB), which caches high-quality training examples with non-zero advantages and reintroduces them during policy updates. This mechanism increases gradient density, enhances training stability, and encourages deeper reasoning paths.
In addition to GRPO, we incorporate a Mixed Preference Optimization (MPO) strategy that fuses preference-based and rule-based reinforcement signals. This hybrid scheme enables a more effective balance between reasoning specialization and generalization. Notably, we find that overly aggressive optimization of reasoning capabilities may lead to increased hallucination in visual tasks, emphasizing the importance of careful reward calibration.
@@
Extensive evaluations demonstrate the effectiveness of R1V2 across multiple authoritative multimodal benchmarks, including 62.6% on OlympiadBench, 79.0% on AIME2024, 63.6% on LiveCodeBench, and 74.0% on MMMU. These results not only establish new open-source baselines but also substantially reduce the performance gap with proprietary state-of-the-art models such as Gemini 2.5 and OpenAI o4-mini.
@@ 2.1 Multimodal Reasoning Models
Recent advances in multimodal AI have increasingly focused on enhancing reasoning capabilities across different modalities. Several approaches have emerged to advance this frontier, broadly categorized into distillation-based methods, reinforcement learning techniques, and specialized architectural innovations.
Distillation-based approaches like Claude-3 [1], GPT-4V [17], and Gemini [23] leverage larger text-only reasoning models to transfer their abilities to multimodal contexts. These models typically use a teacher-student framework where the multimodal model is trained to emulate the reasoning processes of powerful text-only predecessors. While effective, this approach relies heavily on the availability of large proprietary text models, limiting accessibility for open research and creating potential bottlenecks in model diversity and innovation.
Reinforcement learning for multimodal reasoning has gained traction through models like Qwen-VL [3], SPHINX [12], and LLaVA [13]. These methods apply techniques such as RLHF, DPO, and other preference optimization strategies to align model outputs with human preferences for reasoning quality. However, most implementations have primarily focused on general instruction following rather than specialized reasoning capabilities, failing to effectively capture the nuanced progression of complex reasoning chains necessary for mathematical and scientific problem-solving.
Recent architectural innovations include "slow-thinking" approaches like OpenAI-o1 [16], Gemini-Thinking [22], and Kimi-1.5 [24], which introduce specialized mechanisms for extended deliberation. These models allocate additional computational resources to complex reasoning tasks, demonstrating substantial improvements on mathematical and scientific benchmarks. The first generation of Skywork-R1V [19] pioneered the direct application of text reasoning capabilities to vision through advanced adapter techniques, establishing a new paradigm for multimodal reasoning.
However, a common challenge across these approaches is maintaining balance between specialized reasoning abilities and general-purpose multimodal understanding. Models optimized heavily for mathematical reasoning often demonstrate degraded performance on everyday visual tasks, while general-purpose models struggle with complex analytical reasoning. This fundamental tension motivates our current work.
@@ 2.2 Preference Optimization in Multimodal Models
Preference optimization has emerged as a powerful paradigm for aligning AI systems with human expectations, though its extension from language-only to multimodal reasoning presents significant challenges. While foundational techniques like RLHF [5] and DPO [20] have transformed text-based models, their direct application to multimodal contexts requires substantial adaptations to accommodate the complexity of cross-modal interactions.
Early multimodal preference work focused primarily on basic instruction following. Models like BLIP-2 [11] and InstructBLIP [6] pioneered instruction tuning approaches for vision-language alignment, while LLaVA-RLHF [21] and MM-REACT [29] explored human preference incorporation for general multimodal tasks. However, these efforts typically addressed straightforward visual question-answering scenarios rather than complex reasoning chains involving visual evidence.
The specific challenge of preference optimization for multimodal reasoning remains largely unexplored. Although recent models such as DeepSeek-VL [14] and Qwen-VL-Plus [2] have incorporated elements of preference-based training, neither specifically addresses the unique requirements of deep reasoning tasks where visual information must be integrated into extended logical deductions.
Despite these advances, applying preference optimization to complex multimodal reasoning remains significantly underexplored, with two critical limitations hindering progress in this domain. First, the binary nature of typical preference pairs fails to capture the nuanced progression of complex reasoning paths, where multiple equally valid solutions may exist with different intermediate steps. Second, existing reward models predominantly evaluate textual quality in isolation, overlooking the crucial relationship between visual interpretation and logical inference that defines successful multimodal reasoning.
Our work introduces a hybrid optimization framework that transcends these limitations by integrating preference-based signals with structured rule-based reinforcement. This approach provides more granular guidance for reasoning processes—capturing both the quality of intermediate steps and final outcomes—while simultaneously maintaining the model’s general-purpose capabilities across diverse tasks.
@@ 3 Methodology
This section introduces the core methodology behind Skywork R1V2. Building upon the original R1V series, R1V2 incorporates a hybrid training strategy combining rule-based reinforcement learning and reward-model-guided preference learning to better balance reasoning capability and general-purpose performance.
@@ 3.1 Efficient Multimodal Transfer via Modular Reassembly
To reduce reliance on large-scale multimodal reasoning data, we decouple the alignment of visual-language representations from the preservation of reasoning capabilities. Specifically, we use a lightweight multilayer perceptron (MLP) adapter to connect a frozen vision encoder $f_{v}$ (Intern ViT-6B) with a reasoning-capable language model $f_{l}$ (e.g., qwq-32B).
Unlike the first generation, R1V2 omits the supervised fine-tuning stage. Empirical evidence shows that while supervised fine-tuning improves visual-language alignment, it significantly impairs the original reasoning ability of the language model. In contrast, directly combining the pretrained language model with the visual adapter preserves its reasoning ability, albeit with a slight reduction in general visual understanding. This modular approach not only reduces computational requirements but also enables more targeted optimization of cross-modal alignment without disrupting pre-established reasoning capabilities.
@@
3.2 Hybrid Optimization with GRPO and MPO
To balance reasoning and generalization, R1V2 employs Group Relative Policy Optimization (GRPO) and Mixed Preference Optimization (MPO) as its core reinforcement learning strategies.
GRPO: We apply GRPO to guide the model via intra-group reward comparisons between candidate responses. However, we observe a prominent challenge: the "vanishing advantage" problem—when all responses in a group converge to be uniformly correct or incorrect, the relative advantage signal diminishes to near-zero, effectively starving gradient-based policy updates of meaningful learning signals. To address this, we introduce a Selective Sample Buffer (SSB) that stores and reuses high-value training samples with non-zero advantages, improving both training efficiency and stability.
@@ 3.2 Hybrid Optimization with GRPO and MPO
MPO: In parallel with GRPO, we propose a Mixed Preference Optimization strategy that integrates preference signals from the R1V-RM reward model with hand-crafted rule-based constraints (e.g., format correctness, factual consistency, step-by-step reasoning completeness). This hybrid reward structure better aligns the model’s outputs with both stylistic preferences and factual requirements across modalities.
The combination of GRPO and MPO significantly enhances the model’s reasoning and format compliance. After DPO-based fine-tuning, the model outputs become more stable and coherent, with fewer crashes, while retaining a reasonable level of generalization.
We also observe a trade-off between reasoning and general-purpose performance. As a result, our training strategy prioritizes improvements on reasoning benchmarks for the current stage.
3.3 Selective Sample Buffer (SSB) for Addressing Vanishing Advantages
@@
In reinforcement learning approaches, particularly with Group Relative Policy Optimization (GRPO), we observed a significant challenge during training: the “vanishing advantages” problem. This phenomenon occurs when all responses within a query group converge to become uniformly correct or incorrect, causing relative advantage signals to diminish and hindering effective gradient-based policy updates.
@@
Figure 1: The Selective Sample Buffer (SSB) mechanism addresses the vanishing advantages problem by preserving and prioritizing high-value samples with non-zero advantages.
@@
To address this issue, we propose the Selective Sample Buffer (SSB), which operates by identifying and caching high-quality training examples with non-zero advantages from previous iterations. The approach prioritizes samples based on the absolute value of their advantage signals through weighted sampling, strategically reintroducing these informative samples during policy updates. This method maintains a gradient-rich training environment even as model responses converge.
Empirically, we observed that the percentage of effective samples (those with non-zero advantages) decreases dramatically from approximately 60% at the beginning of training to below 10% in later stages, severely impacting training efficiency. The SSB mechanism counteracts this trend by ensuring a consistent supply of informative training signals.
@@
Our experiments demonstrate that the SSB mechanism substantially increases training efficiency, with performance improvements equivalent to full training achieved using just a fraction of the samples. This approach not only addresses the vanishing advantages problem but also prevents the model from plateauing in mid-training, ensuring continuous improvement in reasoning capabilities.

### v2

@@ Abstract
We present Skywork R1V2, a next-generation multimodal reasoning model and a major leap forward from its predecessor, Skywork R1V. At its core, R1V2 introduces a hybrid reinforcement learning paradigm that jointly leverages the Mixed Preference Optimization (MPO) and the Group Relative Policy Optimization (GRPO), which harmonizes reward-model guidance with rule-based strategies, thereby addressing the long-standing challenge of balancing sophisticated reasoning capabilities with broad generalization. To further enhance training efficiency, we introduce the Selective Sample Buffer (SSB) mechanism, which effectively counters the “Vanishing Advantages” dilemma inherent in GRPO by prioritizing high-value samples throughout the optimization process. Notably, we observe that excessive reinforcement signals can induce visual hallucinations–a phenomenon we systematically monitor and mitigate through calibrated reward thresholds throughout the training process. Empirical results affirm the exceptional capability of R1V2, with benchmark-leading performances such as 62.6 on OlympiadBench, 78.9 on AIME2024, 63.6 on LiveCodeBench, and 73.6 on MMMU. These results underscore R1V2’s superiority over existing open-source models and demonstrate significant progress in closing the performance gap with premier proprietary systems, including Gemini 2.5 and OpenAI-o4-mini. The Skywork R1V2 model weights have been publicly released to promote openness and reproducibility***https://huggingface.co/Skywork/Skywork-R1V2-38B.
@@ 1 Introduction
Recent advancements in “slow-thinking” multimodal models—such as OpenAI-o1 [24], Gemini-Thinking [32], Kimi-1.5 [34], and Skywork-R1V [26]—have led to substantial progress in complex reasoning tasks across mathematics and science. These models emulate reflective cognitive processes, achieving stronger performance than their “fast-thinking” counterparts (e.g., GPT-4o [24] and Claude-3.5 [2]), which prioritize rapid generation over deliberate reasoning. For instance, slow-thinking models demonstrate over 30% improvement on benchmarks like AIME24 and AMC23, along with approximately 10% gains on science-related benchmarks such as GPQA [24, 14, 25].
@@
However, extending slow-thinking strategies to multimodal domains introduces new challenges. While improvements are observed on visual reasoning tasks such as MMMU [42] and MathVision [37], performance degrades on general perception benchmarks (e.g., AI2D [18]). This decline is often accompanied by a rise in visual hallucinations, highlighting a critical issue: How can we effectively promote slow-thinking behavior in vision-language models (VLMs) without compromising their generalization capabilities?
@@
To tackle this problem, we introduce Skywork R1V2, a next-generation vision-language model that acquires multimodal reasoning skills directly via reinforcement learning (RL), bypassing the need for teacher model distillation. R1V2 is trained on a comprehensive dataset spanning visual perception, scientific inquiry, and abstract reasoning, encompassing both general-purpose and reasoning-intensive tasks. The training process utilizes a hybrid reinforcement learning paradigm that jointly leverages the Mixed Preference Optimization (MPO) [38] and the Group Relative Policy Optimization (GRPO) [28].
R1V2 first implements MPO with three core learning objectives: (1) the relative preferences between response pairs, (2) the absolute quality of individual responses, and (3) the process for generating preferred responses [38]. The preference signals is provided by our Skywork-VL Reward—a reward model capable of evaluating both multimodal understanding and reasoning tasks. Given these high-quality preference signals, MPO effectively mitigates hallucinations and overthinking in the generated outputs, leading to improved performance on general vision tasks.
@@
To further boost the reasoning capabilities, we incorporate the GRPO algorithm, which computes relative advantages by comparing candidate responses within the same query group. However, as training progresses, candidate responses tend to converge (i.e., become uniformly correct or incorrect), resulting in vanishing advantage signals and limited reasoning diversity. To address this, we employ the Selective Sample Buffer (SSB), which caches high-quality training examples with non-zero advantages and reintroduces them during policy updates. This mechanism increases gradient density, enhances training stability, and encourages deeper reasoning paths. The hybrid of MPO and GRPO enables a more effective balance between reasoning specialization and generalization.
@@
Extensive evaluations demonstrate the effectiveness of R1V2 across multiple authoritative multimodal benchmarks, including 62.6% on OlympiadBench [13], 78.9% on AIME2024, 63.6% on LiveCodeBench [15], and 73.6% on MMMU. These results not only establish new open-source baselines but also substantially reduce the performance gap with proprietary state-of-the-art models such as Gemini 2.5 [8] and OpenAI-o4-mini [25].
@@ 2.1 Multimodal Reasoning Models
Recent advances in multimodal AI have increasingly focused on enhancing reasoning capabilities across different modalities. Proprietary models like Claude-3 [2], GPT-4o [24], and Gemini [33] have demonstrated remarkable capabilities, while the open-source community has responded with competitive alternatives such as LLaVA [21], Qwen-VL [3], and InternVL [6].
Recent innovations have shifted toward slow-thinking approaches (e.g., OpenAI-o1 [24], Gemini-Thinking [32], and Kimi-1.5 [34]), which introduce specialized mechanisms for extended deliberation. These models allocate additional computational resources to complex reasoning tasks, demonstrating substantial improvements on mathematical and scientific benchmarks. The first generation of Skywork-R1V [26] pioneered the direct application of text reasoning capabilities to vision through advanced adapter techniques, establishing a new paradigm for multimodal reasoning.
However, a common challenge across these approaches is maintaining balance between specialized reasoning abilities and general-purpose multimodal understanding. Models optimized heavily for mathematical reasoning often demonstrate degraded performance on everyday visual tasks, while general-purpose models struggle with complex analytical reasoning. Hence, addressing the trade-off between reasoning specialization and generalization is one of the key motivations behind our work.
@@ 2.2 Preference Optimization in Multimodal Models
Preference optimization has emerged as a powerful paradigm for aligning AI systems with human expectations, though its extension from language-only to multimodal reasoning presents significant challenges. While foundational techniques like RLHF [7] and DPO [27] have transformed text-based models, their direct application to multimodal contexts requires substantial adaptations to accommodate the complexity of cross-modal interactions. Recent works on multimodal preference learning [31, 38] have incorporated human preferences into vision-language alignment, enhancing response quality in general multimodal tasks. Despite these advances in addressing straightforward visual question-answering scenarios, the application of preference optimization to complex multimodal reasoning remains relatively underexplored, with two critical limitations hindering progress in this domain. First, the binary nature of typical preference pairs fails to capture the nuanced progression of complex reasoning paths, where multiple equally valid solutions may exist with different intermediate steps. Second, existing reward models predominantly evaluate textual quality in isolation, overlooking the crucial relationship between visual interpretation and logical inference that defines successful multimodal reasoning.
@@ 3 Methodology
This section introduces the core methodology behind Skywork R1V2. Building upon the original R1V series, R1V2 incorporates a hybrid training strategy combining reinforcement learning and reward-model-guided preference learning to better balance reasoning capability and general-purpose performance.
@@ 3.1 Efficient Multimodal Transfer via Modular Reassembly
To reduce reliance on large-scale multimodal reasoning data, we decouple the alignment of visual-language representations from the preservation of reasoning capabilities. Specifically, we introduce a lightweight multi-layer perceptron (MLP) adapter, denote as $f_{c}$, to bridge a frozen vision encoder $f_{v}$ with a reasoning-capable language model $f_{l}$. Here we choose InternViT-6B [6] to be vision encoder and QwQ-32B [36] to be our language model. Formally, given a visual input $x_{v}$ and a text input $x_{t}$, the overall process represented as:
$y=f_{l}(f_{c}(f_{v}(x_{v})),x_{t})$ (1)
where $f_{v}$ extracts visual features, $f_{c}$ adapts these features for capability with the language model, and $f_{l}$ integrates both the adapted visual features and the textual input $x_{t}$ to perform reasoning and generate the output $y$.
Unlike the first generation of R1V, R1V2 eliminates the supervised fine-tuning (SFT) stage. Recent finding suggests that SFT can inadvertently undermine the performance of subsequent reinforcement learning or reasoning processes [5], which can hinder the model’s ability to develop genuine reasoning behaviors. Rather than relying on SFT, R1V2 adopts a modular approach that directly connects a pretrained reasoning language model with a visual adapter. While this approach leads to a slight reduction in general visual understanding, it preserves the inherent reasoning ability of the language model and significantly benefits overall reasoning performance by avoiding the degradation introduced by SFT.
@@
3.2 Mixed Preference Optimization
To achieve a strong model before reinforcement learning and balance reasoning with generalization, we employ Mixed Preference Optimization (MPO) [38], which demonstrates successful in the Internvl series model, as a crucial component of our optimization pipeline. The alignment of our R1V2 model is significantly enhanced by our Skywork-VL reward model which guides our iterative optimization process using MPO. Noticeably, the process yields a substantial reduction in the occurrence of repetitive chain-of-thought (CoT) and overthinking in the generated output.
MPO loss function can be typically expressed as follows:
$\mathcal{L}=w_{1}\mathcal{L}_{\textit{preference}}+w_{2}\mathcal{L}_{\textit{quality}}+w_{3}\mathcal{L}_{\textit{generation}}.$ (2)
$\mathcal{L}_{\textit{preference}}$ is typically a DPO [44] loss which can optimize the relative preference between positive and negative samples:
$\mathcal{L}_{\textit{preference}}=-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{c}\mid x\right)}{\pi_{0}\left(y_{c}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{r}\mid x\right)}{\pi_{0}\left(y_{r}\mid x\right)}\right)$ (3)
Where $\beta$ is the KL penalty coefficient, and the prompt, positive sample, and negative response are represented by $x$, $y_{c}$, and $y_{r}$, respectively. The policy model $\pi_{\theta}$ is initialized from $\pi_{0}$.
$\mathcal{L}_{\textit{quality}}$ is a BCO [17] loss. This loss helps the model to understand the absolute quality of individual responses. This algorithm trains a binary classifier, where the logit serves as a reward and effectively maps the chosen response to 1 and the rejected response to 0. The loss function is defined as:
$\mathcal{L}_{\textit{quality}}$ $=\mathcal{L}_{\textit{quality}}^{+}+\mathcal{L}_{\textit{quality}}^{-}$
$=-\Bigg{[}\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{c}\mid x\right)}{\pi_{0}\left(y_{c}\mid x\right)}-\delta\right)+\log\sigma\left(-\left(\beta\log\frac{\pi_{\theta}\left(y_{r}\mid x\right)}{\pi_{0}\left(y_{r}\mid x\right)}-\delta\right)\right)\Bigg{]}$ (4)
Here, $\delta$ is practically computed as the moving average of past rewards to stabilize training, which is a typically method in online RL method [1, 29].
Furthermore, the generation loss:
$\mathcal{L}_{\textit{generation}}=-\frac{\log\pi_{\theta}(y_{c}|x)}{|y_{c}|}$ (5)
which is typically a Negative Log Likelihood loss (NLL), guiding the model to learn the chosen response, aiming to reduce the distribution shift between the base model’s outputs and the preferred responses.
@@ 3.2 Mixed Preference Optimization
Mixed Preference Optimization strategy can integrate preference signals from the Skywork-VL reward model with hand-crafted rule-based constraints (e.g., format correctness, factual consistency, step-by-step reasoning completeness). This hybrid reward structure better aligns the model’s outputs with both stylistic preferences and factual requirements across modalities.
3.3 Reinforcement Fine-tuning
In the reinforcement fine-tuning stage for VLMs, we primarily employ the GRPO Algorithm with a hybrid supervised signal composed of both rule-based reward and model-based reward. Additionally, we leverage the SSB mechanism to further enhance the efficiency of our reinforcement learning process.
3.3.1 GRPO Algorithm with Hybrid Reward Signal
For enhanced reasoning in multi-modal contexts, we adapt Group Relative Policy Optimization (GRPO) [28], a common RL algorithm originally developed for text-only LLMs. GRPO is a policy optimization algorithm designed to compute token-level advantage estimates by performing intra-group comparisons of generated responses conditioned on a specific query. For a given input instance $x$, a behavior policy $\pi_{\theta_{old}}$ samples a batch of $N$ candidate responses $\{y_{i}\}_{i=1}^{N}$. The advantage $\hat{A}_{i,t}$ for the $i$-th response at time step $t$ is determined by normalizing the rewards obtained across the response group, according to:
$\hat{A}_{i,t}=\frac{r(x,y_{i})-\text{mean}(\{r(x,y_{1}),...,r(x,y_{N})\})}{\text{std}(\{r(x,y_{1}),...,r(x,y_{N})\})}$ (6)
To mitigate the “alignment tax” [20] in reasoning ability, we again utilize our Skywork-VL reward model to introduce a preference reward signal $r_{\theta}$, supplementing the rule-based reward $r_{\textit{rule}}$. Furthermore, we incorporate a format reward $r_{\textit{format}}$ to align the model’s output with the DeepSeek R1-style [9] chat template. Thus, our hybrid reward function is defined as follows:
$r(x,y_{i})=r_{\textit{rule}}(x,y_{i})+r_{\theta}(x,y_{i})+r_{\textit{format}}(x,y_{i})$ (7)
The GRPO optimization objective incorporates a clipped surrogate loss term augmented with a KL-penalty to ensure stable policy updates, formulated as:
$\mathcal{L}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\left[\frac{\pi_{\theta}(y_{i,t}\mid x_{i},y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid x_{i},y_{i,<t})}\hat{A}_{i,t},\ \text{clip}\left(\frac{\pi_{\theta}(y_{i,t}\mid x_{i},y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid x_{i},y_{i,<t})},\ 1-\epsilon,\ 1+\epsilon\right)\hat{A}_{i,t}\right]$ (8)
The hyperparameter $\epsilon$ dictates the tolerance for policy deviation. The clipping function serves to prevent excessively large policy updates by constraining the ratio of the current policy to the reference policy within a predefined interval. This mechanism promotes training stability and mitigates the risk of performance degradation resulting from overly aggressive updates.
3.3.2 Addressing Vanishing Advantages via SSB
@@
However, the direct application of GRPO to VLMs encounters challenges due to its limited effectiveness, particularly after the utilization of MPO. Surprisingly, we observed a detrimental “Vanishing Advantages” phenomenon during GRPO training, which can hinder the upper bound of the optimization. This issue arises when all responses within a query group converge towards uniform correctness or incorrectness, causing the relative advantage signals to diminish and impeding effective gradient-based policy updates.
@@
Figure 1: The Selective Sample Buffer (SSB) mechanism addresses the “Vanishing Advantages” problem by preserving and prioritizing high-value samples with non-zero advantages.
@@
To address this issue, we employ the Selective Sample Buffer (SSB), which operates by identifying and caching high-quality training examples with non-zero advantages from previous iterations. This approach prioritizes samples based on the absolute value of their advantage signals through weighted sampling, strategically reintroducing these informative samples during policy updates. This method maintains a gradient-rich training environment even as model responses converge.
A significant empirical observation is the dramatic reduction in effective training samples (those exhibiting non-zero advantages) from about 60% at the start to under 40% in later phases. This severe drop-off cripples training efficiency. The SSB mechanism actively combats this by guaranteeing a continuous source of valuable training signals. Moreover, we discovered that incorporating SSB into the rollout of offline inference time in advance, effectively creating a filtered prompt pool, yields a substantial improvement of over 10% in training efficiency during the initial optimization.
Our experiments demonstrate that the SSB mechanism substantially increases training efficiency, with performance improvements equivalent to full training achieved using just a fraction of the samples. This approach not only addresses the “Vanishing Advantages” problem but also prevents the model from plateauing in mid-training, ensuring continuous improvement in reasoning capabilities.